OpenStudy (anonymous):

Solve the following system of equations: −2x + y = 1 −4x + y = −1

OpenStudy (waleed_imtiaz):

x=-11 and y=3

OpenStudy (theviper):

-2x+y=-1 & 1-4x+y=-1 =>-4x+y=-1-1=-2

OpenStudy (theviper):

Can u solve it now??

OpenStudy (anonymous):

is it ( -1,3)?

OpenStudy (waleed_imtiaz):

subtract the equations and solve for x an y

OpenStudy (anonymous):

x=-1 y=3

OpenStudy (waleed_imtiaz):

i think so it is (1,3) as I already mentioned.....

OpenStudy (anonymous):

so 1,3?

OpenStudy (waleed_imtiaz):

yes...

OpenStudy (anonymous):

thank you

OpenStudy (waleed_imtiaz):

No problem xD